40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Mon Oct 14 17:01:43 2019, anchal, Notes, Laser, South Laser Switched Off! SouthLaserStopped.pdf
    Reply  Mon Oct 14 18:30:14 2019, anchal, Notes, PMC, South PMC Error Signal Oscillations SPMCMixerOutOscillations.pdf
       Reply  Tue Oct 15 18:54:16 2019, anchal, Notes, PMC, South PMC further problems 
          Reply  Wed Oct 16 16:56:43 2019, anchal, Notes, PMC, South PMC won't work with it's own Servo Card only 
             Reply  Wed Oct 16 17:33:08 2019, Koji, Notes, PMC, South PMC won't work with it's own Servo Card only 
                Reply  Wed Oct 16 18:46:58 2019, anchal, Notes, PMC, South PMC started working somehow :( 
                   Reply  Wed Oct 23 17:39:44 2019, anchal, Notes, PMC, South PMC started working somehow :( 
                      Reply  Fri Oct 25 16:50:12 2019, anchal, Notes, PMC, South PMC problem debugging efforts 
                         Reply  Tue Oct 29 19:10:59 2019, anchal, Notes, PMC, South PMC problem debugging efforts SPMC_Diagnostics.pdf
                            Reply  Wed Oct 30 15:15:28 2019, anchal, Notes, PMC, Replacing U5A AD602 didn't solve the issue. 
                               Reply  Wed Oct 30 18:35:31 2019, anchal, Notes, PMC, Issue closed 
                                  Reply  Mon Nov 4 12:15:01 2019, anchal, Notes, PMC, Changing the autolocking method 
Message ID: 2452     Entry time: Mon Oct 14 18:30:14 2019     In reply to: 2451     Reply to this: 2455
Author: anchal 
Type: Notes 
Category: PMC 
Subject: South PMC Error Signal Oscillations 

South PMC Error signal is showing huge weird capacitor charging type oscillations. Attached is an oscilloscope measurement of it. A more weird thing is that this peak is appearing at random frequencies, that is, if I take single sequence measurements of 1 ms length, I see these peaks occurring at a difference from 100 us to 500 us, randomly.


Changes:

  • I had to re-adjust the power going to the PMC to get 9.6mW power falling on the cavities. But I think the knob was more or less in same place.
  • I had to restart whole lab (all channels and scripts) as I suspected some channels are not responsive.
  • I can't increase the fast gain on southside beyond -7 dB (while this was 14 dB before). so this SPMC oscillations are really hindering badly.
  • I also suspect that the NPRO temperature control is not as stable as before on the south side, probably because it was switched off for more than 2 days.
  • I'm hoping by tomorrow, this should get better. but still need to investigate the issue on south PMC.

Data

Attachment 1: SPMCMixerOutOscillations.pdf  95 kB  Uploaded Mon Oct 14 19:38:16 2019  | Hide | Hide all
SPMCMixerOutOscillations.pdf SPMCMixerOutOscillations.pdf
ELOG V3.1.3-